A non-conducting ring of radius 'r' has charge 'Q'. A magnetic field perpendicular to the plane of th ring changes as dB/dt.the torque experienced by the ring is?

Grade:12

1 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

Dear M.vs vivek vardhan,

 

torque = u*B  (this is cross product)

 

 u = the magnetic dipole moment 

B =the magnetic field vector

u = I*A    = (current * area)

u x B = u*B*Sin(theta)     here theta is 90    and I=dQ/dt

 

Torque=  dQ/dt * B * (Π *r^2)  =   Q* dB/dt  * (Π *r^2)

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay